# Does FLUX Already Know How to Perform Physically Plausible Image Composition?

### v3

@@ 4.2 Experimental Results
We compare our method with two categories of baselines: (1) Training-based methods (6 in total): UniCombine (Wang et al., 2025), AnyDoor (Chen et al., 2024c), Paint by Example (PBE) (Yang et al., 2023), ObjectStitch (Song et al., 2023), MADD (He et al., 2024), and DreamCom (Lu et al., 2023c); (2) Training-free methods (5 in total): EEdit (Yan et al., 2025), TIGIC (Li et al., 2024b), DreamEdit (Li et al., 2023b), TF-ICON (Lu et al., 2023d), and TALE (Pham et al., 2024).
@@ 4.3 Ablation Study
Method
MSA
DSG
ABB
Subject Identity Consistency Background Image Quality
@@ A.1 Image Composition
Training-based methods fine-tune foundational models using datasets tailored for image composition. Early methods like Paint by Example (Yang et al., 2022) and ObjectStitch (Song et al., 2022) use CLIP to encode subject features, ensuring high semantic similarity between inserted objects and reference images. These approaches use image augmentation to create training datasets, enabling effective training. GLIGEN (Li et al., 2023c) incorporates grounding information into new trainable layers of a pre-trained diffusion model via a gated mechanism. ControlCom (Zhang et al., 2023a) integrates 2-dim indicator vector to improve controllability. DreamCom (Lu et al., 2023c) and MureObjectStitch (Chen et al., 2024a) fine-tune models with small sets of reference images to preserve subject identity. AnyDoor (Chen et al., 2024c), IMPRINT (Song et al., 2024b), and E-MD3C (Pham et al., 2025) leverage DINOv2 to enhance identity fidelity and control over shape and pose, drawing supervision from video data. MimicBrush (Chen et al., 2024b) similarly uses video-derived supervision for imitative editing. In contrast, MADD (He et al., 2024), ObjectMate (Winter et al., 2024), and OmniPaint (Yu et al., 2025b) employ image inpainting models to generate higher-quality training datasets compared to those based on image or video augmentation. Multitwine (Tarrés et al., 2025) enables the integration of multiple objects, capturing interactions from simple positional relationships to complex actions requiring reposing. DreamFuse (Huang et al., 2025) uses a Positional Affine mechanism to embed foreground size and position into the background, fostering effective foreground-background interaction through shared attention. Insert Anything (Song et al., 2025) and UniCombine (Wang et al., 2025) introduces a FLUX-based, multi-conditional generative framework that handles diverse condition combinations. However, these methods often bias the generative priors of base models toward curated datasets, resulting in unnatural compositions, such as implausible object-environment interactions (e.g., missing or unrealistic shadows and reflections). This stems from the absence of a large-scale, high-quality, multi-resolution, real-world triplet dataset comprising an object, a scene, and the object seamlessly integrated into the scene, which is expensive to produce.
@@
Training-free methods, on the other hand, modify the inference process of pre-trained models to achieve composition without additional training. Early approaches like TF-ICON (Lu et al., 2023d) leverage accurate image inversion to lay the groundwork for composition, achieved through composite self-attention map injection. TALE (Pham et al., 2024) and PrimeComposer (Wang et al., 2024c) build on TF-ICON to enhance identity preservation and background-object style adaptation. TIGIC (Li et al., 2024b) focuses on preserving non-target areas during composition. Thinking Outside the BBox (Tarrés et al., 2024) enables unconstrained image compositing, unbound by input masks. FreeCompose (Chen et al., 2024e) employs a pipeline of object removal, image harmonization, and semantic composition. DreamEdit (Li et al., 2023b), UniCanvas (Jin et al., 2025), and Magic Insert (Ruiz et al., 2024) use test-time tuning to fine-tune models during inference. Add-it (Tewel et al., 2024) enables text-guided object insertion on FLUX, where users describe objects via text prompts instead of reference images. EEdit (Yan et al., 2025) recently improves TF-ICON on FLUX, introducing step-skipping to reduce time costs and spatial locality caching to minimize redundancy. However, training-free methods rely on precise image inversion and fragile attention surgery, which can lock inserted objects into the exact pose of the reference image, leading to awkward or contextually inappropriate orientations. Attention manipulation often causes instability and hyperparameter sensitivity, as feature or attention map injection does not always preserve subject identity. This creates a trade-off: stronger injection preserves identity but results in unnatural poses, while lighter injection yields more natural poses but compromises identity.
@@ A.2 General Image Editing
Recent and significant advancements in text-to-image generative models have enhanced numerous applications (Zhou et al., 2025c; a; b; 2023; 2024a; 2024b; 2024c; Zhao et al., 2024a; d; c; 2025b; 2025a; b; Xie et al., 2024; Wang et al., 2023b; Wang & Chen, 2025a; b; Lu et al., 2024b; Li et al., 2025; Gao et al., 2024; 2025; Zhu et al., 2024; Yu et al., 2025a). Instruction-based image editing has evolved rapidly. Early systems relied on modular, two-stage pipelines: a multimodal language model first produced textual prompts, spatial guidance, or synthetic instruction–image pairs, and a separate diffusion model then executed the edit—as in InstructEdit (Wang et al., 2023a), InstructPix2Pix (Brooks et al., 2022), MagicBrush (Zhang et al., 2023b), and BrushEdit (Li et al., 2024c). Recent work has shifted toward tightly integrated, instruction-centric architectures. Models such as SmartEdit (Huang et al., 2024), X2I (Ma et al., 2025), RPG (Yang et al., 2024), AnyEdit (Yu et al., 2024), and UltraEdit (Zhao et al., 2024e) embed routing, task-aware objectives, and fine-grained controls directly into the network, yielding higher fidelity and more precise manipulation.
Unified generation-and-editing frameworks (e.g., OmniGen (Xiao et al., 2024), ACE (Han et al., 2024), ACE++ (Mao et al., 2025), Lumina-OmniLV (Pu et al., 2025), Qwen2VL-Flux (Lu, 2024), DreamEngine (Chen et al., 2025), MetaQueries (Pan et al., 2025), Hidream-E1 (HiDream-ai, 2025)) treat editing as one capability of an end-to-end vision-language model, often fusing language embeddings with diffusion latents to provide context-aware, pixel-level control. Efficiency has advanced in parallel: ICEdit (Zhang et al., 2025) couples LoRA with mixture-of-experts tuning and optimized noise initialization, while SuperEdit (Ming et al., 2025) relies on higher-quality data and contrastive supervision to sustain performance at lower cost. Looking ahead, large foundation models such as Gemini (Gemini2.5, 2025) and GPT-5 (OpenAI, 2025) already show strong visual reasoning and coherent, instruction-guided image generation. Yet, despite extensive multimodal pre-training, they still fall short on image composition: object placement remains hard to control, lighting is often inconsistent, and subjects can drift in identity.
@@ A.3 Subject-Driven Generation
Test-time fine-tuning methods (Gal et al., 2022a; Ruiz et al., 2023a) adapt a pre-trained T2I model to a small set of reference images (typically 3 to 5 images). Although this step adds computational cost and latency, it offers the greatest flexibility for diverse customization requirements. Such methods are commonly grouped into three subclasses: data regularization, weight regularization, and loss regularization. In the data-regularization family, DreamBooth (Ruiz et al., 2023a) limits overfitting by generating superclass images with the base T2I model and training on both reference and regularization images; Custom Diffusion (Kumari et al., 2023) improves regularization quality by retrieving real images; and Specialist Diffusion (Lu et al., 2023a) applies extensive data augmentation. Weight-regularization approaches (Gal et al., 2022b; Hu et al., 2021b; Han et al., 2023; Qiu et al., 2023) confine updates to carefully chosen parameters, such as the subject-specific text embeddings or the singular values of weight matrices. Loss-regularization approaches, including Specialist Diffusion (Lu et al., 2023a), MagiCapture (Hyung et al., 2023), and FaceChain-SuDe (Qiao et al., 2024), introduce objective terms that respectively maximize CLIP-space similarity between generated and reference images, disentangle identity and style via masked facial reconstruction, or encourage correct superclass classification.
Zero-shot image customization methods avoid subject-specific fine-tuning at inference time but rely on extensive pre-training. For general subject customization, InstantBooth (Shi et al., 2024a) adds a visual encoder that captures coarse-to-fine image features from the references; BLIP-Diffusion (Li et al., 2024a) fine-tunes BLIP-2 (Li et al., 2023a) to extract multimodal subject representations; ELITE (Wei et al., 2023) maps reference images into hierarchical textual tokens through global and local networks; and Song et al. (Song et al., 2024a) enhance textual control by removing the projection of visual embeddings onto textual embeddings. For facial customization, InstantID (Wang et al., 2024b) isolates facial regions from reference images to extract appearance and structural cues. For style customization, InstantStyle (Wang et al., 2024a) identifies style-controlling layers and injects IP-Adapter features (Ye et al., 2023a) into those layers to achieve style transfer. InstantCharacter (Tao et al., 2025), IP-Adapter (Ye et al., 2023b), and PuLID (Guo et al., 2024) have each released versions compatible with the FLUX model.
@@ Appendix D Evaluating Cross-Attention Map Accuracy via IoU
In the main text, we showcase the effectiveness of our approach on Flux, a leading open-source model. To establish its generalizability, we conduct experiments on SDXL (Podell et al., 2024), SD3.5 (Esser et al., 2024), and PixArt (Chen et al., 2023) across DreamEditBench and ComplexCompo. The results are presented in Tab. 3. On DreamEditBench, our LoRA-based variant with Flux achieves state-of-the-art performance in subject identity preservation, while delivering superior image quality. Similarly, on ComplexCompo, the Flux-LoRA configuration excels in identity consistency and image fidelity. Notably, the framework’s benefits extend beyond a single model family: both SDXL and PixArt-$\Sigma$ exhibit substantial performance gains, affirming the approach’s generality and adaptability across diverse generative architectures.
@@
Bench Method
Base
Model
Subject Identity Consistency Background Image Quality
@@
DreamEdit-
Bench
@@
Complex-
Compo
@@ Appendix G Benchmark Details
Our benchmark consists of 300 triplets, each comprising a subject image, a background image, and a bounding box. The subject images are identical to those used in DreamEditBench (Li et al., 2023b; Ruiz et al., 2023a), while the background images are sampled from the OpenImages dataset (Kuznetsova et al., 2020). These backgrounds exhibit a variety of aspect ratios and resolutions, including landscape and portrait formats, such as $768\times 1088$, $768\times 1072$, $768\times 1024$, $768\times 1152$, $024\times 768, 11$$152\times 768$, $1200\times 768$, $848\times 768$, and $1360\times 768$. The bounding boxes are manually designed to ensure that the size and placement of the inserted subjects are contextually appropriate and visually plausible. The benchmark will be released publicly.
@@ Appendix I Subject Identity Metrics Analysis
In our experiments we found that widely-used subject-identity metrics such as CLIP-I (Radford et al., 2021) and DINOv2 (Oquab et al., 2023) correlate poorly with human preferences. Because they focus almost exclusively on semantic similarity, they ignore appearance changes introduced by lighting, shadows, reflections, and surrounding context. Fig. 12(b) presents several image pairs produced by AnyDoor (left) and by our method (right); the corresponding CLIP-I $(\uparrow)$ (Radford et al., 2021), DINOv2 $(\uparrow)$ (Oquab et al., 2023), IRF $(\uparrow)$ (Shao & Cui, 2022), and DreamSim $(\downarrow)$ (Fu et al., 2023) scores are shown beneath each image, with the better score highlighted in red. Although the AnyDoor results are visibly less realistic and less consistent, they nevertheless receive higher CLIP-I and DINOv2 scores, and in most cases higher IRF scores, demonstrating that these measures do not faithfully capture compositional quality.
@@
The hyperparameters used in our frameworks are summarized in Tab. 5. Each baseline is implemented according to the configuration settings recommended in its original publication. The repositories utilized for each baseline are listed below:
@@
Total steps
$(T-1)\to 0$
Start
step $t_{1}$
Step Range
$t_{1}\to\tau$
Learning Rate
$\alpha$
#iters
$k$
Step Range
$t_{1}\to 0$
Scale
$\eta$
Blur
$\sigma$
Step Range
$(t_{1}-1)\to 1$
Bin
Thresh
Dilation
Kernel Size

### v4

@@ 4.2 Experimental Results
We compare our method with two categories of baselines: (1) Training-based methods (6 in total): UniCombine (Wang et al., 2025a), AnyDoor (Chen et al., 2024c), Paint by Example (PBE) (Yang et al., 2023), ObjectStitch (Song et al., 2023), MADD (He et al., 2024), and DreamCom (Lu et al., 2023c); (2) Training-free methods (5 in total): EEdit (Yan et al., 2025), TIGIC (Li et al., 2024b), DreamEdit (Li et al., 2023b), TF-ICON (Lu et al., 2023d), and TALE (Pham et al., 2024).
@@ 4.3 Ablation Study
Method MSA DSG ABB Subject Identity Consistency Background Image Quality
@@ 5 Conclusion
Ethics Statement
Our framework provides an accessible way for people without professional artistic skills to create image compositions. While this technology offers significant benefits, it also carries the risk of misuse for malicious purposes, such as harassment or spreading misinformation. Additionally, our framework relies on pretrained large-scale T2I models, which may inadvertently introduce social and cultural biases. Therefore, using these models raises ethical concerns and requires careful consideration. We therefore urge users to exercise caution and use this tool responsibly for appropriate purposes only.
Reproducibility Statement
We have taken several measures to ensure the reproducibility of our results. Algorithm 1 presents the pseudocode of our method. Details of the implementation for the main experiments are provided in Sec. 4.1, while the hyperparameter configurations are listed in Appendix J. The source code and the ComplexCompo dataset will be released publicly. This work adheres to the reproducibility requirements set by ICLR.
Acknowledgments
This research is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority.
@@ A.1 Image Composition
Training-based methods fine-tune foundational models using datasets tailored for image composition. Early methods like Paint by Example (Yang et al., 2023) and ObjectStitch (Song et al., 2023) use CLIP to encode subject features, ensuring high semantic similarity between inserted objects and reference images. These approaches use image augmentation to create training datasets, enabling effective training. GLIGEN (Li et al., 2023c) incorporates grounding information into new trainable layers of a pre-trained diffusion model via a gated mechanism. ControlCom (Zhang et al., 2023a) integrates 2-dim indicator vector to improve controllability. DreamCom (Lu et al., 2023c) and MureObjectStitch (Chen et al., 2024a) fine-tune models with small sets of reference images to preserve subject identity. AnyDoor (Chen et al., 2024c), IMPRINT (Song et al., 2024), and E-MD3C (Pham et al., 2025) leverage DINOv2 to enhance identity fidelity and control over shape and pose, drawing supervision from video data. MimicBrush (Chen et al., 2024b) similarly uses video-derived supervision for imitative editing. In contrast, MADD (He et al., 2024), ObjectMate (Winter et al., 2025), and OmniPaint (Yu et al., 2025c) employ image inpainting models to generate higher-quality training datasets compared to those based on image or video augmentation. Multitwine (Tarrés et al., 2025) enables the integration of multiple objects, capturing interactions from simple positional relationships to complex actions requiring reposing. DreamFuse (Huang et al., 2025) uses a Positional Affine mechanism to embed foreground size and position into the background, fostering effective foreground-background interaction through shared attention. Insert Anything (Song et al., 2025a) and UniCombine (Wang et al., 2025a) introduces a FLUX-based, multi-conditional generative framework that handles diverse condition combinations. However, these methods often bias the generative priors of base models toward curated datasets, resulting in unnatural compositions, such as implausible object-environment interactions (e.g., missing or unrealistic shadows and reflections). This stems from the absence of a large-scale, high-quality, multi-resolution, real-world triplet dataset comprising an object, a scene, and the object seamlessly integrated into the scene, which is expensive to produce.
@@
Training-free methods, on the other hand, modify the inference process of pre-trained models to achieve composition without additional training. Early approaches like TF-ICON (Lu et al., 2023d) leverage accurate image inversion to lay the groundwork for composition, achieved through composite self-attention map injection. TALE (Pham et al., 2024) and PrimeComposer (Wang et al., 2024c) build on TF-ICON to enhance identity preservation and background-object style adaptation. TIGIC (Li et al., 2024b) focuses on preserving non-target areas during composition. Thinking Outside the BBox (Canet Tarrés et al., 2024) enables unconstrained image compositing, unbound by input masks. FreeCompose (Chen et al., 2024e) employs a pipeline of object removal, image harmonization, and semantic composition. DreamEdit (Li et al., 2023b), UniCanvas (Jin et al., 2025), and Magic Insert (Ruiz et al., 2025) use test-time tuning to fine-tune models during inference. Add-it (Tewel et al., ) enables text-guided object insertion on FLUX, where users describe objects via text prompts instead of reference images. EEdit (Yan et al., 2025) recently improves TF-ICON on FLUX, introducing step-skipping to reduce time costs and spatial locality caching to minimize redundancy. However, training-free methods rely on precise image inversion and fragile attention surgery, which can lock inserted objects into the exact pose of the reference image, leading to awkward or contextually inappropriate orientations. Attention manipulation often causes instability and hyperparameter sensitivity, as feature or attention map injection does not always preserve subject identity. This creates a trade-off: stronger injection preserves identity but results in unnatural poses, while lighter injection yields more natural poses but compromises identity.
@@ A.2 General Image Editing
Recent and significant advancements in text-to-image generative models have enhanced numerous applications (Zhou et al., 2023; 2024a; 2024b; 2024c; 2025a; 2025b; Wang et al., 2025b; Wang & Chen, 2025a; b; Zhao et al., 2024a; 2025b; 2025a; b; 2026; Li et al., 2025; Gao et al., 2024; 2025; Zhu et al., 2024; Lu et al., 2024b; Ren et al., 2025; Yu et al., 2025b). Instruction-based image editing has evolved rapidly. Early systems relied on modular, two-stage pipelines: a multimodal language model first produced textual prompts, spatial guidance, or synthetic instruction–image pairs, and a separate diffusion model then executed the edit—as in InstructEdit (Wang et al., 2023), InstructPix2Pix (Brooks et al., 2023), MagicBrush (Zhang et al., 2023b), and BrushEdit (Li et al., 2024c). Recent work has shifted toward tightly integrated, instruction-centric architectures. Models such as SmartEdit (Huang et al., 2024), X2I (Ma et al., 2025a), RPG (Yang et al., 2024), AnyEdit (Yu et al., 2025a), and UltraEdit (Zhao et al., 2024c) embed routing, task-aware objectives, and fine-grained controls directly into the network, yielding higher fidelity and more precise manipulation.
Unified generation-and-editing frameworks (e.g., OmniGen (Xiao et al., 2025), ACE (Han et al., ), ACE++ (Mao et al., 2025), Lumina-OmniLV (Pu et al., 2025), Qwen2VL-Flux (Lu, 2024), DreamEngine (Chen et al., 2025), MetaQueries (Pan et al., 2025), Hidream-E1 (HiDream-ai, 2025)) treat editing as one capability of an end-to-end vision-language model, often fusing language embeddings with diffusion latents to provide context-aware, pixel-level control. Efficiency has advanced in parallel: ICEdit (Zhang et al., 2025) couples LoRA with mixture-of-experts tuning and optimized noise initialization, while SuperEdit (Ming et al., 2025) relies on higher-quality data and contrastive supervision to sustain performance at lower cost. Looking ahead, large foundation models such as Gemini (Gemini2.5, 2025) and GPT-5 (OpenAI, 2025) already show strong visual reasoning and coherent, instruction-guided image generation. Yet, despite extensive multimodal pre-training, they still fall short on image composition: object placement remains hard to control, lighting is often inconsistent, and subjects can drift in identity.
@@ A.3 Subject-Driven Generation
Test-time fine-tuning methods (Gal et al., 2022; Ruiz et al., 2023) adapt a pre-trained T2I model to a small set of reference images (typically 3 to 5 images). Although this step adds computational cost and latency, it offers the greatest flexibility for diverse customization requirements. Such methods are commonly grouped into three subclasses: data regularization, weight regularization, and loss regularization. In the data-regularization family, DreamBooth (Ruiz et al., 2023) limits overfitting by generating superclass images with the base T2I model and training on both reference and regularization images; Custom Diffusion (Kumari et al., 2023) improves regularization quality by retrieving real images; and Specialist Diffusion (Lu et al., 2023a) applies extensive data augmentation. Weight-regularization approaches (Gal et al., 2022; Hu et al., ; Han et al., 2023; Qiu et al., 2023) confine updates to carefully chosen parameters, such as the subject-specific text embeddings or the singular values of weight matrices. Loss-regularization approaches, including Specialist Diffusion (Lu et al., 2023a), MagiCapture (Hyung et al., 2024), and FaceChain-SuDe (Qiao et al., 2024), introduce objective terms that respectively maximize CLIP-space similarity between generated and reference images, disentangle identity and style via masked facial reconstruction, or encourage correct superclass classification.
Zero-shot image customization methods avoid subject-specific fine-tuning at inference time but rely on extensive pre-training. For general subject customization, InstantBooth (Shi et al., 2024a) adds a visual encoder that captures coarse-to-fine image features from the references; BLIP-Diffusion (Li et al., 2024a) fine-tunes BLIP-2 (Li et al., 2023a) to extract multimodal subject representations; ELITE (Wei et al., 2023) maps reference images into hierarchical textual tokens through global and local networks; and Song et al. (Song et al., 2025b) enhance textual control by removing the projection of visual embeddings onto textual embeddings. For facial customization, InstantID (Wang et al., 2024b) isolates facial regions from reference images to extract appearance and structural cues. For style customization, InstantStyle (Wang et al., 2024a) identifies style-controlling layers and injects IP-Adapter features (Ye et al., 2023) into those layers to achieve style transfer. InstantCharacter (Tao et al., 2025), IP-Adapter (Ye et al., 2023), and PuLID (Guo et al., 2024) have each released versions compatible with the FLUX model.
@@ C.2 Blurring the Key and Value Matrices
C.3 Implementation Details of Gaussian Blurring
For the 2D Gaussian filtering step, we adopt the implementation provided by the kornia library. Following standard engineering practice, we set the kernel radius to $r=3\sigma$, since three standard deviations capture approximately $99.7\%$ of the Gaussian mass. Consequently, the kernel size is chosen as the nearest odd integer to $2r=6\sigma$. In all of our experiments we use $\sigma=10$ (see Tab. 5).
The procedure is applied to query embeddings by reshaping them into spatial feature maps, performing Gaussian smoothing, and then mapping them back into the sequence domain prior to attention computation. The full workflow is given below:
Algorithm 2 Implementation Details of 2D Gaussian Blurring
1:Input: A query matrix $\mathbf{Q}\in\mathbb{R}^{B\times L\times D}$, key $\mathbf{K}$, value $\mathbf{V}$, spatial dimensions $(H,W)$, Gaussian standard deviation $\sigma$.
2:Output: Attention output $\mathbf{O}$.
3:// Reshape query into spatial tensor
[…]
6:// Construct Gaussian kernel size
[…]
11:// Apply 2D Gaussian smoothing
[…]
13:// Reshape smoothed queries back to sequence form
[…]
16:// Compute attention
[…]
19:return $\mathbf{O}$
@@ Appendix D Evaluating Cross-Attention Map Accuracy via IoU
In the main text, we showcase the effectiveness of our approach on Flux, a leading open-source model. To establish its generalizability, we conduct experiments on SDXL (Podell et al., 2024), SD3.5 (Esser et al., 2024), and PixArt (Chen et al., ) across DreamEditBench and ComplexCompo. The results are presented in Tab. 3. On DreamEditBench, our LoRA-based variant with Flux achieves state-of-the-art performance in subject identity preservation, while delivering superior image quality. Similarly, on ComplexCompo, the Flux-LoRA configuration excels in identity consistency and image fidelity. Notably, the framework’s benefits extend beyond a single model family: both SDXL and PixArt-$\Sigma$ exhibit substantial performance gains, affirming the approach’s generality and adaptability across diverse generative architectures.
@@
Bench Method Base Model Subject Identity Consistency Background Image Quality
@@
DreamEdit- Bench (220) Flux.1 Fill FLUX 0.7328 0.6745 0.5754 0.5233 0.0166 0.9076 0.5577 3.5997
@@
Complex- Compo (300) Flux.1 Fill FLUX 0.7108 0.6475 0.5466 0.6018 0.0232 0.7442 0.4088 3.5737
@@ Appendix G Benchmark Details
Our benchmark consists of 300 triplets, each comprising a subject image, a background image, and a bounding box. The subject images are identical to those used in DreamEditBench (Li et al., 2023b; Ruiz et al., 2023), while the background images are sampled from the OpenImages dataset (Kuznetsova et al., 2020). These backgrounds exhibit a variety of aspect ratios and resolutions, including landscape and portrait formats, such as $768\times 1088$, $768\times 1072$, $768\times 1024$, $768\times 1152$, $024\times 768, 11$$152\times 768$, $1200\times 768$, $848\times 768$, and $1360\times 768$. The bounding boxes are manually designed to ensure that the size and placement of the inserted subjects are contextually appropriate and visually plausible. The benchmark will be released publicly.
@@ Appendix I Subject Identity Metrics Analysis
In our experiments we found that widely-used subject-identity metrics such as CLIP-I (Radford et al., 2021) and DINOv2 (Oquab et al., 2024) correlate poorly with human preferences. Because they focus almost exclusively on semantic similarity, they ignore appearance changes introduced by lighting, shadows, reflections, and surrounding context. Fig. 12(b) presents several image pairs produced by AnyDoor (left) and by our method (right); the corresponding CLIP-I $(\uparrow)$ (Radford et al., 2021), DINOv2 $(\uparrow)$ (Oquab et al., 2024), IRF $(\uparrow)$ (Shao & Cui, 2022), and DreamSim $(\downarrow)$ (Fu et al., 2023) scores are shown beneath each image, with the better score highlighted in red. Although the AnyDoor results are visibly less realistic and less consistent, they nevertheless receive higher CLIP-I and DINOv2 scores, and in most cases higher IRF scores, demonstrating that these measures do not faithfully capture compositional quality.
@@
The hyperparameters used in our frameworks are summarized in Tab. 5. Under Denoising Setup, “Total steps” refers to the full diffusion/noising schedule, which specifies a sequence of 20 values of $\sigma_{t}$ across timesteps $t$. However, our method does not begin denoising at the first timestep. As shown in Algorithm 1, it starts at $t_{1}=14$ (Ours-Adapter), resulting in 15 denoising steps in total. Under MSA loss, “Step Range” indicates the subset of denoising steps to which MSA optimization is applied. For the adapter setting, the MSA loss is applied only to the first three denoising steps (from $t=14$ to $t=12$). In addition, since a LoRA is trained for a specific subject, it provides a more tailored prior than the generic adapter, allowing it to converge in fewer steps and with lower overall compute.
Each baseline is implemented according to the configuration settings recommended in its original publication. The repositories utilized for each baseline are listed below:
@@
Total steps $(T-1)\to 0$ Start step $t_{1}$ Step Range $t_{1}\to\tau$ Learning Rate $\alpha$ #iters $k$ Step Range $t_{1}\to 0$ Scale $\eta$ Blur $\sigma$ Step Range $(t_{1}-1)\to 1$ Bin Thresh Dilation Kernel Size